# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ 7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$. The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines $1$ to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
The key difference is that the total number of jobs $n$ is unknown in advance. At any time $t$, we only observe the number of jobs released so far, denoted by $n(t)$, rather than the final value of $n$. Consequently, the rank-based partition can no longer use a fixed threshold; instead, it must classify jobs as large or small using a dynamic threshold $\ell(t)$ that depends on $n(t)$. This introduces a major difficulty: the partition is no longer stable as in the known-$n$ setting.
Previously, when $n$ was known, the threshold $\ell$ was fixed. A small job would remain small forever, and the only source of instability was when a large job retired, which we handled by introducing a proxy job. In contrast, in the unknown-$n$ setting, $\ell(t)$ increases monotonically as $n(t)$ grows, and thus varies over time.
As a result, a job may switch between being large and small multiple times. For instance, when $n(t)$ is small, we may treat the top $3$ jobs as large, so a job of rank $4$ is small. Later, as more jobs arrive and $\ell(t)$ becomes $4$, this job becomes large. If even more jobs arrive, $\ell(t)$ may increase to $5$, but that same job may drop out of the top-$5$ and become small again. This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
To this end, we make two main modifications as follow in the new algorithm as follows, and the formal description is presented in Algorithm 5.
More Dynamic Partition Rule.
We use a partition rule that depends on the number of jobs released so far. Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use $N$ to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$. Let $T_{j}(t)$ denote the type of job $j$ at time $t$. We classify $T_{j}(t)$ as large if job $j$ is among the top $\ell$ jobs in size and $p_{j}\geq\frac{4P(t)}{\ell}$, where $P(t)$ is the total size of all jobs released by time $t$. Otherwise, $T_{j}(t)$ is classified as small.
A New Kill Event When $N$ Is Updated.
Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine. However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier. To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever $N$ is updated (that is, when $n(t)>N$, and we double $N$ to $2N$). At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so. We remark that this is one reason for using a rounded value of $n(t)$ (the nearest power of two) as the benchmark, since it allows us to bound the number of such kill events.
Data: Estimate of $n$: NN initiated as 11;
@@
On any job completion or after any job release at $t$:
@@
If the total waiting time after $t^{\prime}$ is already bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$, then $\sum_{j\in J}F^{(1)}_{j}$ is also bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$.
@@
  1. 1.
    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
  2. 2.
    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the running jobs will be restarted later.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts, $F_{j}^{(A)}$ and $F_{j}^{(B)}$, corresponding to genuine and fake busy intervals:
$t_{j}-t^{\prime}_{j}=F_{j}^{(A)}+F_{j}^{(B)}.$
Bounding $F_{j}^{(A)}$.
@@
$\sum_{j\in J}F_{j}^{(A)}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}\right).$ (9)
Bounding $F_{j}^{(B)}$.
Now consider fake busy intervals caused by type-B kill events. Let the $i$-th type-B kill event occur at time $\beta_{i}$. For a time $t$ to contribute to $F_{j}^{(B)}$, it must satisfy:
  1. 1.
    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
  2. 2.
    $t$ is not a genuine busy time.
We claim that
$\beta_{i}-t\leq p_{j}.$ (10)
Proof of the claim. Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Assume for contradiction that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, so their processing times are at least $p_{j}$. By the algorithm’s rule, job $j$ cannot have been released before starting any of these larger jobs. Let $U$ denote this set of running jobs. Then
@@
However, by definition,
$t^{\prime}_{j}=r_{j}+\frac{(2n)^{\alpha}}{\sqrt{m}}\cdot p_{j},$
and let $P(r_{j})$ denote the total workload released up to $r_{j}$, then
$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i},$
contradicting $\beta_{i}<t^{\prime}_{j}$. Thus, (10) holds.
Each kill event therefore contributes at most $\sum_{j\in J}p_{j}\leq\textup{{OPT}}$ to $\sum_{j}F_{j}^{(B)}$. Since the total number of type-B kill events is at most $O(n^{1-\alpha}/\sqrt{m})$, we obtain
$\sum_{j\in J}F_{j}^{(B)}\leq O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$ (11)
@@
$\sum_{j\in J}F^{(1)}_{j}\leq\sum_{j\in J}\bigl((t^{\prime}_{j}-r_{j})+F_{j}^{(A)}+F_{j}^{(B)}\bigr)\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$
@@
    Blocked: $j\in S_{h}$, and some machine is processing a job $j^{\prime}$ at $\theta_{j}$, where $\theta_{j}<\theta_{j^{\prime}}$. We use $B_{h}$ to denote the set of blocked small jobs in $S_{h}$.
@@
We focus on the number of waiting blocked jobs at time $t^{*}$, which is at most $n^{\alpha}\sqrt{m}$. Let $B^{*}_{h}$ denote this set. View $t^{*}$ as time 0 and form two shifted instance $U^{\prime}_{h}$ and $B^{\prime}_{h}$ by moving every release time in $U_{h}$ and $B^{*}_{h}$ by $t^{*}$, repectively. The schedule of $U_{h}$ in our algorithm is a shifting to the schedule of $U^{\prime}_{h}$ under $\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b})$, where $\vec{b}$ is the initial blocking vector from the jobs processing at $t^{*}$. By the definition of $t^{*}$, these processing jobs are already in their last small period, so each component of $\vec{b}$ is at most $O\!\left(\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\right)$.
@@
$\sum_{j\in U_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(U^{\prime}_{h}\cup B^{\prime}_{h},\vec{b}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}},$ (12)
@@
The next step is to bound the gap between $\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h})$ and $\textup{{OPT}}^{m}(U_{h})$ using the fact that $|B^{\prime}_{h}|$ is small. Consider an optimal preemptive, migratory schedule $\textup{{OPT}}^{m}(U^{\prime}_{h})$. Insert the jobs of $B^{\prime}_{h}$ one by one with migration: for each $j\in B^{\prime}_{h}$, we process it in the earliest idle slots without delaying already scheduled work; if the machine is busy, pause and resume at the next idle slot (possibly on another machine).
@@
$F^{\prime}_{j}\;=\;p_{j}+\text{wait}_{j},\qquad\text{wait}_{j}\;\leq\;\frac{1}{m}\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
By the lemma’s condition, $|B^{\prime}_{h}|=|B^{*}_{h}|\leq n^{\alpha}\sqrt{m}$, hence
$F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}.$
@@
$\sum_{j\in U_{h}}F^{(2)}_{j}$ $\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup B^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup B^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
$\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup B^{\prime}_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
@@
$\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}\;\leq\;O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
@@
Case 2. Some Small-Only machines are not busy at the end. This only happens at a type-A kill event. If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment. Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when $N$ was updated to $2N$ — the only time a small job can become large. At such an update, at most $N$ jobs have been released, so there are at most $N$ blocked jobs. Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$. Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$ at most
@@
There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 5, that is $O(n^{\alpha}/\sqrt{m})$-competitive against the preemptive offline solution without the prior knowledge of $n$, where $\alpha=(\sqrt{5}-1)/2$, provided that $m\geq 2$.
@@ 8.2 Generalization to Multiple Machines
Moving to the analysis of an arbitrary online algorithm, we aim to prove that the algorithm’s flow time is at least $\Omega(mk^{2})$. First, we will show that when a large job is placed such that it conflicts with the small jobs (having a intersection of at least $0.5$), it significantly increases the flow time of those small jobs.
@@
Any randomized online algorithm for minimizing the total flow time on parallel machines has a competitive ratio of at least $\Gamma(n,m)$, for every fixed pair $(n,m)$, where $\Gamma(n,m)=\Omega(\sqrt{n/m})$.
@@
$\mathbb{E}(F^{(L)})$ $\geq\sum_{i=0}^{k-1}\left((i+1)\cdot m-\sum_{j=0}^{i-1}\mathbb{E}(N_{b})\right)\geq\mathbb{E}\left(\sum_{i=0}^{k-1}\sum_{b=0}^{i}\left(m-\mathbb{E}(N_{b})\right)\right)$
@@
By Lemma 21 $\textup{{OPT}}=O(mk)$, any deterministic algorithm is then $\Omega(k)=\Omega(\sqrt{n/m})$ competitive on the counterexample, and we can conclude the theorem by Yao’s minimax principle. ∎

### v3

@@ 7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
In this section, we extend the $O(\sqrt{n/m})$-competitive kill-and-restart algorithm of Section 6 to the setting where the total number of jobs $n$ is unknown in advance. The resulting algorithm achieves a competitive ratio of $O(n^{\alpha}/\sqrt{m})$, where $\alpha=(\sqrt{5}-1)/2$.
The overall framework is the same: we keep the machine partition into Small-Only and Mixed machines, and we still kill large jobs once enough small jobs are blocked. The key challenge is in the partition. Recall that in the known-$n$ setting, the threshold $\ell$ was fixed, making the partition nearly stable: a small job could never become large (since $\ell$ does not grow), and the only reclassification was in the opposite direction — a large job could be retired and handled via a proxy job.
Without knowledge of $n$, the threshold $\ell(t)$ must depend on the number of jobs $n(t)$ released so far, and it increases over time. This means a job’s classification can oscillate: for example, a job of rank $4$ is small when $\ell=3$, becomes large when $\ell$ grows to $4$, and may become small again when $\ell$ reaches $5$ and larger jobs take its place. This instability invalidates the previous analysis.
To address this, we make two modifications, described below and formalized in Algorithm 5.
Dynamic partition rule.
We round $n(t)$ up to the nearest power of two, $N=2^{\lceil\log_{2}n(t)\rceil}$, and set the threshold $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$. At every time $t$, a job $j$ is classified as large if it is among the top $\ell$ jobs by processing time and $p_{j}\geq 4P(t)/\ell$, where $P(t)$ is the total processing time of all jobs released by time $t$; otherwise, it is classified as small. The rounding ensures that $\ell$ changes only $O(\log n)$ times.
Additional kill event on threshold update.
When $N$ doubles, some jobs currently running on Small-Only machines may now be classified as large (since $\ell$ has increased and new jobs may have displaced them in rank). We kill all such jobs immediately. This is a new type of kill event, in addition to the existing one triggered when too many small jobs are blocked.
Data: Estimate of $n$: NN initialized to 11;
@@
On any job completion or after any job release:
@@
If the total waiting time after the shifted times $t^{\prime}_{j}$ is already bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$, then $\sum_{j\in J}F^{(1)}_{j}$ is also bounded by $O(n^{\alpha}/\sqrt{m})\textup{{OPT}}$.
@@
  1. $(\checkmark)$
    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the jobs running on Mixed machines are not restarted later.
  2. $(\times)$
    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the jobs running on Mixed machines will be restarted later.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(\checkmark)}$ and $F_{j}^{(\times)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
$t_{j}-t^{\prime}_{j}=F_{j}^{(\checkmark)}+F_{j}^{(\times)}.$
Bounding $F_{j}^{(\checkmark)}$ (genuine busy).
@@ 2. $(\times)$
$\sum_{j\in J}F_{j}^{(\checkmark)}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\cdot\textup{{OPT}}\right).$ (8)
Bounding $F_{j}^{(\times)}$ (fake busy).
Now consider fake busy intervals. These must be caused by type-B kill events, since type-A kill events only kill jobs on Small-Only machines. Let the $i$-th type-B kill event occur at time $\beta_{i}$. The key observation is the following.
Claim 2.
For any $t\in[t’_{j},t_{j})$ that is a fake busy time between consecutive type-B kill events $\beta_{i-1}$ and $\beta_{i}$, we have $\beta_{i}-t\leq p_{j}$.
Proof of 2.
Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Suppose, for contradiction, that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$. By our algorithm, job $j$ had not yet been released when the algorithm started one of these larger jobs. Therefore, letting $U$ denote this set of jobs, we have
@@
However, by the partition rule $p_{j}\geq 4P(r_{j})/\ell$ and $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$, we have $t’_{j}=r_{j}+(2n)^{\alpha}/\sqrt{m}\cdot p_{j}\geq r_{j}+4P(r_{j})/m$. Since the jobs in $U$ were released before $r_{j}$, their total size is at most $P(r_{j})$. Moreover, $|U|>\lceil m/4\rceil$, so:
$\frac{4\sum_{u\in U}p_{u}}{m}\geq\frac{4\cdot|U|\cdot\min_{u\in U}p_{u}}{m}>\min_{u\in U}p_{u}.$
Combining, $t’_{j}\geq r_{j}+4P(r_{j})/m\geq r_{j}+4\sum_{u\in U}p_{u}/m>r_{j}+\min_{u\in U}p_{u}>\beta_{i}$. This gives $\beta_{i}-t\leq\beta_{i}-t’_{j}<0$, a contradiction. ∎
Each kill event therefore contributes at most $\sum_{j\in J}p_{j}\leq\textup{{OPT}}$ to $\sum_{j}F_{j}^{(\times)}$. Since the total number of type-B kill events is at most $O(n^{1-\alpha}/\sqrt{m})$, we obtain
$\sum_{j\in J}F_{j}^{(\times)}\leq O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$ (9)
@@
$\sum_{j\in J}F^{(1)}_{j}\leq\sum_{j\in J}\bigl((t^{\prime}_{j}-r_{j})+F_{j}^{(\checkmark)}+F_{j}^{(\times)}\bigr)\leq O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\cdot\textup{{OPT}}.$
@@
    Blocked: $j\in S_{h}$, and some machine is processing a job $j^{\prime}$ at $\theta_{j}$, where $\theta_{j}<\theta_{j^{\prime}}$. We use $K_{h}$ to denote the set of blocked small jobs in $S_{h}$.
@@
We focus on the number of waiting blocked jobs at time $t^{*}$, which is at most $n^{\alpha}\sqrt{m}$. Let $K^{*}_{h}$ denote this set. View $t^{*}$ as time 0 and form two shifted instances $U^{\prime}_{h}$ and $K^{\prime}_{h}$ by moving every release time in $U_{h}$ and $K^{*}_{h}$ by $t^{*}$, respectively. The schedule of $U_{h}$ in our algorithm is a shifting to the schedule of $U^{\prime}_{h}$ under $\textup{{NSJF}}(U^{\prime}_{h}\cup K^{\prime}_{h},\vec{b})$, where $\vec{b}$ is the initial blocking vector from the jobs processing at $t^{*}$. By the definition of $t^{*}$, these processing jobs are already in their last small period, so each component of $\vec{b}$ is at most $O\!\left(\frac{\textup{{OPT}}}{n^{\alpha}\sqrt{m}}\right)$.
@@
$\sum_{j\in U_{h}}F^{(2)}_{j}\;\leq\;F(\textup{{NSJF}}(U^{\prime}_{h}\cup K^{\prime}_{h},\vec{b}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup K^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}},$ (10)
@@
The next step is to bound the gap between $\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h})$ and $\textup{{OPT}}^{m}(U_{h})$ using the fact that $|K^{\prime}_{h}|$ is small. Consider an optimal preemptive, migratory schedule $\textup{{OPT}}^{m}(U^{\prime}_{h})$. Insert the jobs of $K^{\prime}_{h}$ one by one with migration: for each $j\in K^{\prime}_{h}$, we process it in the earliest idle slots without delaying already scheduled work; if the machine is busy, pause and resume at the next idle slot (possibly on another machine).
@@
$F^{\prime}_{j}\;=\;p_{j}+\text{wait}_{j},\qquad\text{wait}_{j}\;\leq\;\frac{1}{m}\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}.$
By the lemma’s condition, $|K^{\prime}_{h}|=|K^{*}_{h}|\leq n^{\alpha}\sqrt{m}$, hence
$F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}.$
@@
$\sum_{j\in U_{h}}F^{(2)}_{j}$ $\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}\cup K^{\prime}_{h}))\;+\;O\!\left(\frac{|U^{\prime}_{h}\cup K^{\prime}_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
$\leq\;F(\textup{{OPT}}^{m}(U^{\prime}_{h}))\;+\;O\!\left(\frac{n^{\alpha}}{\sqrt{m}}\right)\sum_{i\in U^{\prime}_{h}\cup K^{\prime}_{h}}p_{i}\;+\;O\!\left(\frac{|S_{h}|}{n^{\alpha}\sqrt{m}}\right)\textup{{OPT}}$
@@
$\sum_{h\in H_{2}}\sum_{j\in K_{h}}F^{(2)}_{j}\;\leq\;O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right)\textup{{OPT}}.$
@@
Case 2. Some Small-Only machines are not busy at the end. This only happens at a type-A kill event. If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment. Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when $N$ was updated to $2N$ — the only time a small job can become large. At such an update, at most $N$ jobs have been released, so there are at most $N$ blocked jobs. Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$. Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in K_{h}}F^{(2)}_{j}$ at most
@@
Algorithm 5 is an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization that is $O(n^{\alpha}/\sqrt{m})$-competitive against the preemptive offline solution without prior knowledge of $n$, where $\alpha=(\sqrt{5}-1)/2$, provided that $m\geq 2$.
@@ 8.2 Generalization to Multiple Machines
Moving to the analysis of an arbitrary online algorithm, we aim to prove that the algorithm’s flow time is at least $\Omega(mk^{2})$. First, we will show that when a large job is placed such that it conflicts with the small jobs (having an intersection of at least $0.5$), it significantly increases the flow time of those small jobs.
@@
Any randomized online algorithm for total flow time minimization on parallel machines has a competitive ratio of at least $\Omega(\sqrt{n/m})$, for every fixed pair $(n,m)$.
@@
$\mathbb{E}(F^{(L)})$ $\geq\sum_{i=0}^{k-1}\left((i+1)\cdot m-\sum_{b=0}^{i-1}\mathbb{E}(N_{b})\right)\geq\mathbb{E}\left(\sum_{i=0}^{k-1}\sum_{b=0}^{i}\left(m-\mathbb{E}(N_{b})\right)\right)$
@@
By Lemma 22, $\textup{{OPT}}=O(mk)$, then any deterministic algorithm is $\Omega(k)=\Omega(\sqrt{n/m})$ competitive on the counterexample, and we can conclude the theorem by Yao’s minimax principle. ∎